# Dynamic Reflections: Probing Video Representations with Text Alignment

### v1

@@ 1 Introduction
In recent years, significant progress has been made towards this goal. The success of multimodal learning approaches has highlighted the potential of joint training paradigms (Radford et al., 2021; Yin et al., 2024; Liu et al., 2024b; Dubey et al., 2024; Team et al., 2023; 2025), while a parallel line of research has provided evidence that even unimodal models, when trained at scale, develop representations with strong structural similarities (Huh et al., 2024; Maniparambil et al., 2024a; Tjandrasuwita et al., 2025). This observation has given rise to the Platonic Representation Hypothesis (PRH), which posits that neural networks converge towards a shared statistical model of reality in their latent spaces (Huh et al., 2024). Building on this, several recent works have established a wide range of techniques to evaluate and exploit cross-modal alignment, in both zero-shot and few-shot scenarios (Maniparambil et al., 2024a; Jha et al., 2025; Schnaus et al., 2025; Hadgi et al., 2025). However, prior work has focused almost exclusively on the alignment of static modalities: primarily images and text. In other words, although the Platonic Representation Hypothesis has been stated in full generality, so far, its validity has only been evaluated on static data, leaving its applicability to dynamic, temporal domains an open question. Consequently, the rich information contained in video: motion, causality and temporal dependencies, exhibited, e.g., in intricate human interactions (Gu et al., 2018; Wang et al., 2023), has been largely overlooked in the context of representation alignment.
A notable limitation of previous studies focusing on static modalities was raised in Huh et al. (2024), where the authors pointed out: “the maximum theoretical value for the alignment metric is 1. Is a score of 0.16 indicative of strong alignment […] or does it signify poor alignment with major differences left to explain? We leave this as an open question.” In this work, we provide a partial answer to this open question by showing that the limited alignment observed previously is, in large part, due to the impoverished data given at test time. Specifically, we demonstrate that by considering multiple video frames instead of a single image, as well as a diverse set of captions instead of a single annotation, the alignment score can be improved significantly, achieving close to 0.4 in some cases, without modifying the underlying trained models. This result highlights, for the first time, that large improvements in alignment can be achieved through efforts at test time, which is complementary to the training-time resources (model size, amount of training data, etc.) considered in prior work. It also suggests that multi-frame approaches can capture different aspects of a scene better, and points to the possibility of a strong zero-shot evaluation metric capable of probing the representation power of both image and video encoders.
@@ 6 Video-Text Alignment and Data Dependence
This parametric analysis resembles compute-optimal scaling laws, such as those identified for pre-training, e.g., in Hoffmann et al. (2022), which model the dependency of a loss function (e.g., Negative Log-Likelihood) on compute or data size. Whereas those laws are formulated in terms of a loss metric (where lower is better), our metric is an alignment score (where higher is better), making our additive saturation model conceptually equivalent. The saturation score $S_{\infty}$ (e.g., $0.41$ for VideoMAEv2) can be interpreted as the base accuracy of an ideal alignment process, while the subtracted terms represent the respective error penalties incurred by the approximation of providing only a finite data at inference time.
@@ 9 Conclusion, Limitations, and Future Work
In this work, we conduct the first comprehensive study extending the Platonic Representation Hypothesis into the temporal domain. We demonstrate that alignment scores dramatically improve – doubling in some cases – simply by utilizing multiple video frames and diverse caption sets at inference, without any retraining. We quantify this phenomenon with a novel, highly accurate ($R^{2}>0.98$) saturation-based scaling law, which quantitatively confirms that native video models (like VideoMAEv2) are fundamentally more able to leverage temporal information than static encoders (like DINOv2).
@@ A.3 Synthetizing captions for PVD dataset using LLMs
Human caption  The video shows a person making a burger. The person is adding mayonnaise to one bun and then placing different vegetables and a patty on top of it after that applies ketchup on another bun. The person is also using wooden boards to keep the ingredients and the burger. The person is wearing a red and white checkered shirt.
@@
Figure 9: Example of video frames, the original human-provided caption, and LLM-synthesized captions for a video in the PVD dataset.

### v2

@@ 1 Introduction
In recent years, significant progress has been made towards this goal. The success of multimodal learning approaches has highlighted the potential of joint training paradigms (Radford et al., 2021; Yin et al., 2024; Liu et al., 2024b; Dubey et al., 2024; Team et al., 2023; 2025), while a parallel line of research has provided evidence that even unimodal models, when trained at scale, develop representations with strong structural similarities (Huh et al., 2024; Maniparambil et al., 2024a; Tjandrasuwita et al., 2025). This observation has given rise to the Platonic Representation Hypothesis (PRH), which posits that neural networks converge towards a shared statistical model of reality in their latent spaces (Huh et al., 2024). Building on this, several recent works have established a wide range of techniques to evaluate and exploit cross-modal alignment, in both zero-shot and few-shot scenarios (Maniparambil et al., 2024a; Jha et al., 2025; Schnaus et al., 2025; Hadgi et al., 2025).
However, prior work has focused almost exclusively on the alignment of static modalities: primarily images and text. In other words, although the Platonic Representation Hypothesis has been stated in full generality, so far, its validity has only been evaluated on static data, leaving its applicability to dynamic, temporal domains an open question. Consequently, the rich information contained in video: motion, causality and temporal dependencies, exhibited, e.g., in intricate human interactions (Gu et al., 2018; Wang et al., 2023), has been largely overlooked in the context of representation alignment. A notable limitation of previous studies focusing on static modalities was raised in Huh et al. (2024), where the authors pointed out: “the maximum theoretical value for the alignment metric is 1. Is a score of 0.16 indicative of strong alignment […] or does it signify poor alignment with major differences left to explain? We leave this as an open question.” In this work, we provide a partial answer to this open question by showing that the limited alignment observed previously is, in large part, due to the impoverished data given at test time. Specifically, we demonstrate that by considering multiple video frames instead of a single image, as well as a diverse set of captions instead of a single annotation, the alignment score can be improved significantly, achieving close to 0.4 in some cases, without modifying the underlying trained models. This result highlights, for the first time, that large improvements in alignment can be achieved through efforts at test time, which is complementary to the training-time resources (model size, amount of training data, etc.) considered in prior work. It also suggests that multi-frame approaches can capture different aspects of a scene better, and points to the possibility of a strong zero-shot evaluation metric capable of probing the representation power of both image and video encoders.
@@ 6 Video-Text Alignment and Data Dependence
This parametric analysis resembles compute-optimal scaling laws, such as those identified for pre-training, e.g., in Hoffmann et al. (2022), which model the dependency of a loss function (e.g., Negative Log-Likelihood) on compute or data size. Whereas those laws are formulated in terms of a loss metric (where lower is better), our metric is an alignment score (where higher is better), making our additive saturation model conceptually equivalent. The saturation score $S_{\infty}$ (e.g., $0.41$ for VideoMAEv2) can be interpreted as the base accuracy of an ideal alignment process, while the subtracted terms are the respective error penalties incurred by having a finite approximation of the data at inference time. We provide additional justification for scaling laws in Appendix, Section B.
@@ 9 Conclusion, Limitations, and Future Work
In this work, we conduct the first comprehensive study extending the Platonic Representation Hypothesis into the temporal domain. We demonstrate that alignment scores dramatically improve – doubling in some cases – simply by utilizing multiple video frames and diverse caption sets at inference, without any retraining. We quantify this phenomenon with a novel and accurate ($R^{2}>0.98$) saturation-based scaling law, which quantitatively confirms that native video models (like VideoMAEv2) are intrinsically better at using temporal information than static encoders (like DINOv2).
@@ Appendix A Appendix
  • Section B: Theoretical justification for scaling laws. We justify our framework of scaling laws from first principles using two separate theoretical models drawn from literature.
@@ A.3 Synthetizing captions for PVD dataset using LLMs
Detailed caption  The person is adding mayonnaise and vegetables to the burger. The person is also using a wooden board to make the burger. The video is a tutorial on how to make a burger. The person is wearing a red and white checkered shirt. The video is well lit and the colors are vibrant. The video is a great tutorial for people who want to learn how to make a burger.
@@
Figure 9: Example of video frames, the original fine-grained model caption, and LLM-synthesized captions for a video in the PVD dataset.
@@ A.5 Cross-model (pairwise video-video) alignment)
Appendix B Theoretical Justification for Test-time Scaling Laws
B.1 Introduction
Our paper observes a close relationship between the amount of data provided at test time and the resulting cross-modal alignment score between vision and text encoders, as measured by the Mutual $k$-NN metric introduced in Huh et al. (2024). This relationship is captured by the following parametric model:
$\text{score}(n_{f},n_{c})=S_{\infty}-(C_{f}n_{f}^{-\alpha}+C_{c}n_{c}^{-\beta})$ (3)
This formulation achieves remarkably high coefficients of determination ($R^{2}>0.98$). In the paper, we state this relationship as an empirical observation and call it a “test-time scaling” law, taking inspiration from empirical training-time scaling laws, e.g., Hoffmann et al. (2022).
In this document, our goal is to provide a theoretical justification for these scaling laws. For this, we proceed in the three major stages:
  1. 1.
    First, we provide an interpretation of the observed scaling laws, in terms of different constants involved, and their conceptual meaning.
  2. 2.
    We then provide a justification for the particular form of the scaling law (why subtractive, why power-law, etc.).
  3. 3.
    Lastly, we deduce the empirical test-time scaling laws from theoretical models of how different encoders process data.
B.2 Stage 1: Interpretation of the Scaling Law and its Constants
The scaling law in Eq. equation 3 models the alignment score as a saturation process. As test-time data increases, performance approaches an asymptotic limit. The constants characterize the inherent limitations of the vision and language encoders and the efficiency with which they integrate information.
  1. 1.
    Saturation Score ($S_{\infty}$): $S_{\infty}$ represents the theoretical maximum alignment score achievable by the specific pair of vision and text encoders, assuming access to infinite test-time information ($n_{f}\to\infty,n_{c}\to\infty$). This constant reflects the structural similarity of the latent spaces learned during training, related to the concept of “Platonic alignment” (Huh et al., 2024). It is analogous to the irreducible error in standard learning theory (e.g., Hastie et al. (2009) Chap. 2.9).
  2. 2.
    Reducible Error: The term $(C_{f}n_{f}^{-\alpha}+C_{c}n_{c}^{-\beta})$ represents the “alignment gap.” This is the penalty incurred due to the impoverished information provided by a finite number of samples at inference time.
    1. (a)
      Error Coefficients: ($C_{f}$ and $C_{c}$) These constants quantify the magnitude of the reducible error attributable to the visual ($C_{f}$) and textual ($C_{c}$) modalities, respectively. A larger $C_{f}$ suggests the model has a greater capacity to utilize information from that modality (e.g., VideoMAEv2’s utilization of temporal information compared to DINOv2). Another way of looking at these coefficients, is that they capture the dependence of approaching the irreducible error, using information from a particular modality. Note that when given one caption and one frame, our test-time scaling law becomes: $S_{\infty}-C_{f}-C_{c}.$
    2. (b)
      Scaling Exponents: ($\alpha$ and $\beta$) These exponents determine the rate at which the alignment score approaches $S_{\infty}$. They reflect how efficiently the encoders integrate new information. A larger exponent implies faster saturation. For example, our finding that DINOv2 (image model) has a higher $\alpha$ than VideoMAEv2 (video model) suggests that static models quickly extract relevant visual features but gain little from subsequent frames, whereas native video models integrate temporal information more effectively over time.
B.3 Stage 2: Justification for the Form of the Scaling Law
The specific mathematical form of Equation 3—saturation, subtractive error terms, and power-law dependence—is consistent with established principles in statistical learning and the broader literature on neural network scaling laws (Hestness et al., 2017; Kaplan et al., 2020).
B.3.1 Why Subtractive Form (Saturation)?
The formulation $\text{Performance}=\text{Max Potential}-\text{Error}$ is necessary because the alignment score (Mutual $k$-NN) is increasing (higher is better) and bounded. As the aggregated representation becomes richer, the error must decrease, and the score saturates at $S_{\infty}$. This structure is directly analogous to the scaling laws observed for training dynamics, notably the Chinchilla laws (Hoffmann et al., 2022), which model loss $L$ as a function of model size $N$ and dataset size $D$:
$L(N,D)=E+\frac{A}{N^{a}}+\frac{B}{D^{b}}$ (4)
Here, $E$ is the irreducible error. Since the alignment score is an accuracy metric (higher is better) rather than a loss metric (lower is better), the subtractive form in our Equation 3 is the appropriate adaptation.
B.3.2 Why Power Law?
The power-law relationship ($n^{-\alpha}$) between error and resources (data size, model size, or, here, test-time samples) is ubiquitous in deep learning (Kaplan et al., 2020; Bahri et al., 2021). This arises due to several interconnected factors:
  1. 1.
    Diminishing Returns and Redundancy: In natural data (videos and language), samples are highly correlated. Initial frames or captions capture the most salient aspects, while subsequent samples often provide redundant information. A power law effectively captures this smooth decay in marginal information gain (Bahri et al., 2021; Cohen & Xu, 2015).
  2. 2.
    Statistical Structure of Natural Data: Natural signals often exhibit heavy-tailed distributions of feature importance (e.g., Zipf’s Law) (Zipf, 2016; Newman, 2005). Power-law scaling naturally emerges when learning from such distributions, as shown in Section B.4 below.
  3. 3.
    Data Manifold Geometry: The complexity of the data manifold dictates how many samples are required to approximate it accurately. Approximation theory suggests error scales as a power of the sample size, inversely related to the intrinsic dimensionality (Sharma & Kaplan, 2020).
B.3.3 Why Additive Error Terms ($C_{f}+C_{c}$)?
Our model assumes the total error is the sum of the error from limited frames and the error from limited captions: $E_{total}=E_{f}(n_{f})+E_{c}(n_{c})$. This implies that the information gained from adding more frames (e.g., capturing dynamics) is largely independent of the information gained from adding more captions (e.g., semantic diversity), as these two sources of information are captured by two different (vision and language) encoders. The remarkably high $R^{2}$ values reported in the paper (>0.98>0.98) validate that this additive model is an excellent empirical approximation.
B.4 Stage 3: Deduction from Theoretical Models
In addition to the interpretation provided above, we deduce the observed scaling laws from theoretical models of how encoders acquire and process information. We note that our observed power-law behavior is closely related to “resolution-limited” scaling described in previous scaling law analyses (Bahri et al., 2021). In this regime, performance is limited by how well the available resources (here, test-time samples) can resolve the details of the underlying data structure. We present two complementary frameworks:
B.4.1 Statistical Estimation Setup
  1. 1.
    Idealized Representations and Alignment: For a given event, we assume there exist idealized, “true” representations in the vision ($\bm{V}^{*}$) and text ($\bm{T}^{*}$) embedding spaces. In the context of alignment, we assume that ($\bm{V}^{*}$) and text ($\bm{T}^{*}$) represent the idealized alignable latent subspaces of the learned latent spaces of the two encoders, i.e., $\bm{V}^{*}=\bm{T}^{*}$.
  2. 2.
    Noisy Estimation from Finite Samples: The representations produced by the two encoders from a finite number of samples are noisy estimates of the ideal representation:
    $\bm{V}(n_{f})$ $=\bm{V}^{*}+\bm{\eta}_{V}(n_{f})$ (5)
    $\bm{T}(n_{c})$ $=\bm{T}^{*}+\bm{\eta}_{T}(n_{c})$ (6)
    Here, $\bm{\eta}_{V}$ and $\bm{\eta}_{T}$ are zero-mean noise vectors ($\mathbb{E}[\bm{\eta}]=0$) representing the estimation error due to finite sampling.
  3. 3.
    Independence of Noise Processes: We assume that the noise process in the vision modality ($\bm{\eta}_{V}$) is statistically independent of the noise process in the text modality ($\bm{\eta}_{T}$). This is plausible as the sources of variation (e.g., specific temporal sampling of frames or visual occlusion vs. linguistic ambiguity or phrasing variations in captions) are distinct.
  4. 4.
    Alignment Metric and Mean Squared Error (MSE): We assume the alignment error $E(n_{f},n_{c})$ (the gap between the best possible alignment for the two encoders $S_{\infty}$ and the achieved score) is proportional to the Mean Squared Error (MSE) between the noisy estimates. While the empirical metric is Mutual k-NN, MSE provides a tractable analytical proxy for representation distance.
    $E(n_{f},n_{c})=S_{\infty}-\text{Score}(n_{f},n_{c})\propto\mathbb{E}\left[\|\bm{V}(n_{f})-\bm{T}(n_{c})\|^{2}\right]$ (7)
  5. 5.
    Submanifold Hypothesis: Following previous literature on scaling law analysis, we assume that the data corresponding to an event (e.g., the temporal evolution of the video), as perceived by a particular encoder lies on a $d$-dimensional submanifold within the high-dimensional embedding space (Bengio et al., 2013). This encoder-specific dimensionality reflects the variability in the structure perceived by the encoder. Test-time data ($n$ frames or captions) provides samples from this submanifold. The trained encoder aggregates these samples to estimate the precise representation or the local structure of the manifold corresponding to that specific event.
Deduction Part 1: Derivation of Additivity
The additive structure of the error terms follows from the independence of the noise processes. Specifically, decomposing the MSE and using the definition of the ideal alignment ($\bm{V}^{*}=\bm{T}^{*}$), we have:
$\mathbb{E}\left[\|\bm{V}(n_{f})-\bm{T}(n_{c})\|^{2}\right]$ $=\mathbb{E}\left[\|(\bm{V}^{*}+\bm{\eta}_{V})-(\bm{T}^{*}+\bm{\eta}_{T})\|^{2}\right]$ (8)
$=\mathbb{E}\left[\|\bm{\eta}_{V}\|^{2}+\|\bm{\eta}_{T}\|^{2}-2\bm{\eta}_{V}^{T}\bm{\eta}_{T}\right]$ (9)
$=\mathbb{E}\left[\|\bm{\eta}_{V}\|^{2}\right]+\mathbb{E}\left[\|\bm{\eta}_{T}\|^{2}\right]-2\mathbb{E}\left[\bm{\eta}_{V}^{T}\bm{\eta}_{T}\right].$ (10)
Since $\bm{\eta}_{V}$ and $\bm{\eta}_{T}$ are assumed to be independent and zero-mean, the expectation of the cross-term is zero: $\mathbb{E}[\bm{\eta}_{V}^{T}\bm{\eta}_{T}]=\mathbb{E}[\bm{\eta}_{V}]^{T}\mathbb{E}[\bm{\eta}_{T}]=0$. Therefore, the total alignment error is the sum of the variances of the estimation errors in each modality:
$\mathbb{E}\left[\|\bm{V}(n_{f})-\bm{T}(n_{c})\|^{2}\right]=\underbrace{\mathbb{E}\left[\|\bm{\eta}_{V}(n_{f})\|^{2}\right]}_{E_{V}(n_{f})}+\underbrace{\mathbb{E}\left[\|\bm{\eta}_{T}(n_{c})\|^{2}\right]}_{E_{T}(n_{c})}.$ (11)
Deduction Part 2: Deriving the Power Law via Manifold Approximation
According to approximation theory on manifolds, the error in estimating the structure of the manifold from $n$ samples typically scales as $n^{-1/d}$, where the proportionality constant is related to the smoothness of the manifold or the specific task (see  Levina & Bickel (2004), and Theorems 2 and 3 in Bahri et al. (2021) relating this behavior explicitly to neural network scaling laws). When using the Mean Squared Error (an analytic loss function minimized at the target), the error scales quadratically with the distance. As detailed in Section 2.2.3 of Bahri et al. (2021), the estimation variance (population loss) scales as $E(n)\propto(\Delta x)^{2}\propto n^{-2/d}$. Applying this result to the variance terms in our decomposition yields:
$E_{V}(n_{f})\propto n_{f}^{-2/d_{v}}\quad\text{and}\quad E_{T}(n_{c})\propto n_{c}^{-2/d_{t}},$ (12)
where $d_{v}$ and $d_{t}$ are the effective dimensions as perceived by the vision and text encoders respectively. Substituting these back into our definition of the alignment metric, Eq. equation 7, and the additive error model, Eq. equation 11, we obtain the final scaling law:
$\text{Score}(n_{f},n_{c})=S_{\infty}-\left(C_{f}n_{f}^{-\alpha}+C_{c}n_{c}^{-\beta}\right)$ (13)
where the scaling exponents are given by $\alpha=2/d_{v}$ and $\beta=2/d_{t}$.
Final Form and Interpretation:
Combining the additive structure derived from noise independence (Eq. 11) and the power-law variance decay, we recover the empirical scaling law:
$\text{Score}(n_{f},n_{c})=S_{\infty}-(C_{f}n_{f}^{-\alpha}+C_{c}n_{c}^{-\beta}).$ (14)
  • Image Models (e.g., DINOv2): Image models largely ignore temporal relationships. The effective temporal dimensionality they perceive ($d$) is low. This results in a large $\alpha$ (fast saturation), as additional frames quickly become redundant for approximating the static scene.
  • Video Models (e.g., VideoMAEv2): Video models capture spatiotemporal features, perceiving a higher effective temporal dimensionality ($d$). This results in a smaller $\alpha$ (slower saturation), as more frames are needed to map the temporal dynamics.
Another way to interpret this model is that the scaling exponents directly relate to the redundancy of the information source as perceived by the encoder. For example, image models (e.g., DINOv2) primarily capture static features; additional frames quickly become correlated regarding static content, leading to a relatively large $\alpha$ (faster saturation). Native video models (e.g., VideoMAEv2) capture dynamic, spatiotemporal features where information unfolds over time; this results in a smaller $\alpha$ (slower saturation), as more frames are required to reduce the variance related to temporal dynamics.
B.4.2 Model 2: Modality-Specific Feature Dynamics
We also provide an alternative probabilistic model based on the statistics of natural data.
Assumptions:
  1. 1.
    Modality-Specific Zipfian Distributions: In this model, we assume a countable universe of features indexed by $k$. Furthermore, the importance (weight) of these features depends on the modality, and follows a Zipfian distribution, characteristic of natural data (Cancho & Solé, 2003):
    • Vision Modality: Features have weights $w^{(v)}_{k}$ following a Zipfian distribution $w^{(v)}\propto k^{-\mu_{v}}$.
    • Text Modality: Features have weights $w^{(t)}_{k}$ following a Zipfian distribution $w^{(t)}\propto k^{-\mu_{t}}$.
    Crucially, the ranking of features may differ between modalities (e.g., a visually salient background color may have high $w^{(v)}$ but low $w^{(t)}$).
  2. 2.
    Independent Acquisition Priority: Encoders capture features in the order of importance within their modality. Given $n_{f}$ frames, the video encoder captures the set of features $K_{V}(n_{f})$ corresponding to the top-ranked visual features. We assume that the effective number of features captured in $n_{f}$ frames scales as $|K_{V}|\propto n_{f}$. Similarly, given $n_{c}$ captions, the text encoder captures the set $K_{T}(n_{c})$ corresponding to the top-ranked semantic features, scaling as $|K_{T}|\propto n_{c}$. Note that this is analogous to assuming $D$ data points resolve the top $D$ eigenfunctions in kernel analysis (Bahri et al., 2021).
  3. 3.
    Alignment Definition: The alignment score $A(n_{f},n_{c})$ is defined as the total weight of features captured by both encoders, given $n_{f}$ frames and $n_{c}$ captions. Because alignment reflects the joint information capacity, the contribution of a matched feature $k$ is a function of its weights in both modalities, $\Omega_{k}(w^{(v)}_{k},w^{(t)}_{k})$. For the alignment error to be bounded by the tail of the distributions, we assume the total potential alignment $S_{\infty}=\sum_{k}\Omega_{k}$ is finite.
Deduction:
The total Alignment error $E(n_{f},n_{c})$ represents the potential alignment mass that was missed because either the video or text encoder failed to capture the necessary features. Formally:
$E(n_{f},n_{c})=S_{\infty}-\sum_{k\in K_{V}\cap K_{T}}\Omega_{k}=\sum_{k\notin K_{V}\cap K_{T}}\Omega_{k}$ (15)
Using the Principle of Inclusion-Exclusion, the set of missed features is the union of features missed by Vision ($M_{V}=K_{V}^{c}$) and features missed by Text ($M_{T}=K_{T}^{c}$). Thus:
$E(n_{f},n_{c})=\underbrace{\sum_{k\in M_{V}}\Omega_{k}}_{E_{vid}(n_{f})}+\underbrace{\sum_{k\in M_{T}}\Omega_{k}}_{E_{text}(n_{c})}-\underbrace{\sum_{k\in M_{V}\cap M_{T}}\Omega_{k}}_{E_{overlap}}$ (16)
Now, consider the vision error term $E_{vid}(n_{f})$. This represents the alignment potential lost because the video encoder did not capture the tail features. Even if we assume only a loose correlation between visual and text weights, the total value of features in the “visual tail” is dominated by the power-law decay of the visual weights $w^{(v)}$. We approximate this tail sum using an integral, which is standard practice for analyzing the asymptotics of heavy-tailed distributions (formalized by the Euler-Maclaurin formula) and is commonly used in scaling law derivations (e.g., Bahri et al. (2021), Sec 2.3.3):
$E_{vid}(n_{f})\approx\int_{n_{f}}^{\infty}x^{-\mu_{v}}dx\propto n_{f}^{-(\mu_{v}-1)}$ (17)
This yields the power law $C_{f}n_{f}^{-\alpha}$ where $\alpha=(\mu_{v}-1)$. Applying the same logic, the text error term follows $E_{text}(n_{c})\approx C_{c}n_{c}^{-\beta}$ where $\beta=(\mu_{c}-1)$.
The interaction term $E_{overlap}$ in Eq. 16 represents the mass of features missed by both encoders. In the asymptotic regime (large $n$), we assume that the probability of a feature being in the deep tail of both distributions simultaneously is of lower order than being in the tail of just one. Thus, $E_{overlap}\ll E_{vid}+E_{text}$, and the additive approximation holds:
$Score(n_{f},n_{c})\approx S_{\infty}-(C_{f}n_{f}^{-\alpha}+C_{c}n_{c}^{-\beta}).$ (18)
Final Form and Discussion:
Since $E(n_{f},n_{c})=S_{\infty}-A(n_{f},n_{c})$, substituting the expression of $E(n_{f},n_{c})$ obtained above, we get: $A(n_{f},n_{c})\approx S_{\infty}-(C_{f}n_{f}^{-\alpha}+C_{c}n_{c}^{-\beta}),$ which replicates the empirical scaling law described in our paper.
Observe that there were several explicit assumptions made above, both in relation to the independence of feature acquisition processes, and associated with the negligeable weight corresponding to the interaction term $E_{overlap}$. While these assumptions are consistent with prior work, a more careful investigation would shed better light on the exact ideal form (potentially involving a more careful model of the interactions between modalities, as suggested by this model) as interesting future work.
B.5 Conclusion
All three models provide justification for the observed test-time scaling laws. They demonstrate that the power-law form arises naturally from the statistical properties of natural data (heavy tails) and the geometric properties of the representation space (intrinsic dimensionality), consistent with the resolution-limited framework (Bahri et al., 2021). These models also allow us to connect the empirical parameters ($S_{\infty},\alpha,\beta$) to the fundamental capabilities of the encoders processing the data.